Poiseuille’s law asserts that the speed of blood that is r centimeters from the central axis of an artery of radius S(r)=c(R^2-r^2), where c is a positive constant. Where is the speed of the blood greatest?

1
Expert's answer
2021-11-30T09:44:03-0500

We consider the function "S(r)=c\cdot(R^2-r^2)" on the segment "r\in[0,R]" and find its derivative "S'(r)=(c\cdot(R^2-r^2))_r^{'}=-2c\cdot r\le0,t\in[0,R]" and "S'(r)<0.r\in (0,R]" . Therefore the function S(r) strictly decreases in [0,R] and we have inequality "S(0)>S(r) \forall(r\in(0.R]" . Therefore "S(0)=c\cdot(R^2-0^2)=c\cdot R^2=max_{r\in[0,R]}S(r)" .

So speed of the blood has its maximum as r=0 or in the central axis of an artery.